# ACC Saturator: Automatic Kernel Optimization for Directive-Based GPU Code

### v1

@@ Abstract.
In this paper, we propose equality saturation to optimize sequential codes utilized in directive/based programming for GPUs. Our approach simultaneously realizes less computation, less memory access, and high memory throughput. Our fully/automated framework constructs single/assignment forms from inputs to be entirely rewritten while keeping dependencies and extracts optimal cases. Through practical benchmarks, we demonstrate a significant performance improvement on several compilers. Furthermore, we highlight the advantages of computational reordering and emphasize the significance of memory/access order for modern GPUs.
1. Introduction
@@
Over the last decade, the increasing popularity of computational accelerators in high-performance computing (HPC) has significantly impacted the programming landscape. The emergence of complex, unique, yet highly efficient architectures has posed a challenge for scientists, who must exert additional effort to take advantage of these architectures. Among the fastest computers, Graphics Processing Units (GPUs) are the most pervasive accelerators (TOP500.org, 2022) and are inherently parallel, requiring software to expose concurrent calculations. Traditional optimization techniques may not provide performance benefits in such scenarios. Furthermore, with the approach of the end of Moore’s Law (Moore, 1998), accelerators have become increasingly domain-specific, and compilers have become essential for the automatic utilization of supercomputing resources.
@@
Although coding with low-level languages is typically the most profitable approach, utilizing them often involves complex data dependencies. As a result, adapting vendor/specific languages like CUDA (NVIDIA Corporation, 2022) and OpenCL (The Khronos Group Inc, 2023) requires extensive knowledge of the application and creates separate code entities just for computation offloading, which increases programming costs. To address this challenge, current GPU accelerators support programming with abstract models. OpenACC (Organization, 2011) and OpenMP (ARB, 1997) have emerged as popular models that alleviate the difficulty of accelerator use by providing code directives to existing languages. These directives allow users to specify compute/intensive parts of the original code, which compilers may automatically translate into accelerator code. At runtime, offloading is realized without any user intervention.
@@
Directive-based code typically maintains a general programming style and remains unspecialized until compilation. It is the responsibility of compilers to handle the computation, the redundancy of memory accesses, and the order of those while considering the nature of accelerators. However, code rewriting while preserving semantics may be challenging. Current compilers apply optimization for each metric sequentially, which limits opportunities for improved performance. In particular, user-specified directives fix loop structures, so little work has been done on optimising code under directives other than relying on compiler techniques used for sequential programs (Barua et al., 2018; Matsumura et al., 2021). As a result, more complex approaches are needed to improve the performance of directive-based code.
@@
This paper proposes the use of a modern optimization technique, equality saturation (Tate et al., 2009), to fine-tune directive/based GPU code. Our tool, ACC Saturator444The source code will be available upon publication., achieves less computation, fewer memory accesses, and high throughput at the same time, while easily integrating into the compilation of both OpenACC and OpenMP compilers. ACC Saturator performs kernel optimization by passing programs through an e-graph, a graph structure for equality saturation. Unlike other optimization methods, our approach neither transforms the abstract syntax trees nor it changes directives. Despite this, we attain significant performance improvements of up to 2.23x with the NVHPC compiler and 5.08x with GCC. We present a detailed performance analysis that highlights the benefits of each optimization on the state-of-the-art GPU architecture, NVIDIA Tesla A100. Our contributions are four-fold:
@@
The rest of the paper is organized as follows. Section 2 provides background knowledge on GPUs, directive-based programming, code optimization, and equality saturation. Section 3 presents a high-level overview of our work. Sections 4 and 5 explain our approach to represent general sequential code for equality saturation and optimizations. Section 6 describes our novel code generation technique from e-graphs towards high/throughput GPU execution. In Section 7, we describe our experimental methodology. Section 8 presents the results of our evaluation along with in-depth analysis on each technique. Section 9 discusses related work. Section 10 provides concluding remarks.
2. Background
@@ 2.1. GPUs
Parallel computing is an assemblage of sequential execution, with means of intercommunication and mutual effects on efficiency. Such architectures accommodate simultaneous operations on multiple different data in uniform processors that are adjacent to each other. As an instance, a Graphics Processing Unit (GPU) runs hundreds of thousands of threads at the same time on around a hundred of streaming multiprocessors (SM). One of the most recent GPUs, NVIDIA Tesla A100, features 108 SMs and each SM contains 32 FP64 cores along with multiple LD/ST units (NVIDIA Corporation, 2020).
@@
The primary benefit of GPU execution stems from overlapping utilization of both computational and memory units. Although GPUs provide high memory bandwidth, the global memory access latency is also higher than that of CPUs (Jia et al., 2018, 2019; Velten et al., 2022). Therefore, optimal throughput may be attained by covering memory requests with computational execution and hiding the latency of data movement. Additionally, the order of memory accesses is essential due to the memory hierarchy, because the distance among accesses by neighboring threads often results in limited bandwidth (Zohouri and Matsuoka, 2019).
2.2. Directive-Based Programming
@@
Decomposing sequential execution into parallel entities is challenging. Routine calls often create data dependencies across multiple files, hindering the extraction of compute-intensive code. Even with extracted code, loops often possess intricate access patterns on different structures that challenge automatic parallelization for efficiency. Directive-based programming models, such as OpenACC (Organization, 2011) and OpenMP (ARB, 1997), alleviate the burden of introducing additional code for hardware acceleration. Both OpenACC and OpenMP provide code directives on the de facto standard languages for scientific applications (C/C++/ Fortran) to specify offloading parts as compute kernels with the declaration of explicit parallelism, enabling seamless automatic accelerator use.
@@
Listing 1 provides an example of OpenACC code that utilizes code directives to specify the parallelism of loops. The kernels directive guards a sequence of kernels, while the loop directive explicitly sets the parallelism of a loop. Users may also specify a single kernel region by using the parallel directive (not shown in the listing). OpenACC’s parallelism consists of three levels of abstraction: gang, worker and vector (in coarse to fine order). For instance, on the NVHPC compiler, both the top i and the middle j loops feature gang parallelism with the degree of the trip count cy and 16, respectively, and these are distributed over thread-blocks on GPUs. The middle j loop exposes additional vector parallelism and launches 256 threads to execute the inner statements. The blue comments show an equivalent form in OpenMP, which does not allow the reuse of parallelism across nested loops. Loop parallelism and data transfers may be implicit, and the compiler automatically solves data dependencies and sets optimal parameters. The NVHPC compiler generates embarrassingly parallel code for GPUs from directive code. On the other hand, GCC utilizes a principal-agent model for both OpenACC and OpenMP, and Clang follows the same approach for OpenMP (Matsumura et al., 2023).
2.3. Equality Saturation
@@
Compiler techniques have been critical in utilizing the best resources available for program execution on the target architecture for the last half-century (Aho et al., 2006; Wolfe et al., 1995). Various features of hardware, including registers, cache systems, pipelines, and parallelization, rely on the compiler’s efforts for performance, enabling target-specific optimization under resource constraints. However, compilers tend to miss a holistic view of performance opportunities and require additional discrete efforts for each discrete optimization mechanism (Rawat et al., 2018). Equality saturation, a state/of/the/art technique for compilers (Tate et al., 2009), defines a set of rewriting rules over the accumulation of equal expressions of target code. The rewriting continues until the expression gets saturated, finding no other forms, or hits a time or size limitation, attaining optimal codes based on a cost model that considers the entire computation.
@@
A dedicated graph structure called e-graph accepts the accumulation of equal expressions, while sharing redundant code blocks over expressions. The graph structure consists of e-classes, groups of equal e-nodes. One e-node may point to e-classes as its subexpressions, and the accumulation works by extending or merging e-classes in accordance with rewriting rules, while preserving the relationship between parental e-nodes and child e-classes (Willsey et al., 2021; de Moura and Bjørner, 2007). To extract optimal solutions, each e-class selects one contained e-node with a minimum cost, while optionally counting common expressions only once and completing common subexpression elimination (CSE).
3. Methodology Overview
@@
[…]
4. Program Representation
During compiler optimizations, the program semantics and behavior are preserved while improving its performance. Maintaining the order of data accesses is especially important to ensure reproducibility of results. To express clear dependencies among statements, program analysis often relies on static single-assignment form (SSA), which allows only one definition per variable (Appel, 1997). Early work on equality saturation (Tate et al., 2009) used an SSA-based graph structure to represent and rewrite an entire program. For directive/based code, however, user-specified parallelism significantly affects performance (Montgomery et al., 2015), so compilers are limited to respect users’ decisions.
@@ 4. Program Representation
4.1. E-Graph Creation
@@
Both OpenACC and OpenMP contain sequential parts within the innermost parallel loops that may be executed independently across multiple threads and thus optimized for more efficient execution or reduced computation. To attain this, we create an e-graph for each innermost parallel loop as follows: First, we introduce conditional nodes (Havlak, 1994) to represent control structures such as if and for, while merging data flows. Second, we assign an ID to each variable/array assignment or . Third, we update each variable/array load to refer to the latest ID along its data flow. Last, for each assignment or , we assign both the ID and the expression to the same e-class.
[…]
4.2. Code Selection
@@
To select optimal codes in the e-graph, we aim to find an equal expression that corresponds to a set of all assignments. We extract the lowest-cost expression that contains all the e-classes of assignments based on a sequence of IDs. The total cost is calculated as the sum of the cost of each e-class, with common e-classes being counted only once. To attain this, we use linear programming techniques (Forrest and Lougee-Heimer, 2005).
@@
Each assignment is updated based on its ID and the extracted expression. To enable reuse across assignments, the values of common expressions are stored in temporary variables. ACC Saturator allows for customizing the generation order of these temporary variables (Section 6).
5. Optimization with Saturation
In this section, we present our approach to rewriting OpenACC/OpenMP code using equality saturation. Our strategy involves a carefully crafted set of rewriting rules and a cost model that aims to minimize both computation and memory access.
5.1. Rewriting Rules
[…]
@@ 5.1. Rewriting Rules
We apply two sets of rewriting rules to our equality saturation process. The first set introduces fused multiply/adds (FMA) operations, which can improve code generation and resource utilization on GPUs (Rawat et al., 2018). Table 1 lists the minimum set of rules we use. When we encounter expressions that match the FMA pattern, we add the corresponding operations to the e-classes of matched expressions. The second set of rules benefits from the commutative and associative properties of the plus and multiply operators to reorder computation. This can enable common subexpression elimination and produce new FMA operations. We also incorporate constant folding of arithmetic operations with integer and floating-point numbers.
@@
5.2. Cost Model
GPUs are complex systems that require careful analysis to accurately predict their efficiency (Khairy et al., 2020; Wu et al., 2015). While applications running on GPUs often face memory-bound limitations (Vijaykumar et al., 2015), it is not just the number of memory accesses that affects overall bandwidth. Factors such as on-chip resource utilization, processor occupancy, thread/ grid-level parallelism across multiple memory layers, and instruction-level parallelism (ILP) all play a role, and improving one metric often comes at the expense of another. In ACC Saturator, our focus is on reducing memory access and computation by utilizing registers for common expressions, while maintaining ILP as described in Section 6.
@@ 6. Code Generation
6.1. Temporary-Variable Insertion
[…]

### v3

@@ Abstract
In this paper, we propose equality saturation to optimize sequential codes utilized in directive/based programming for GPUs. Our approach realizes less computation, less memory access, and high memory throughput simultaneously. Our fully/automated framework constructs single/assignment forms from inputs to be entirely rewritten while keeping dependencies and extracts optimal cases. Through practical benchmarks, we demonstrate a significant performance improvement on several compilers. Furthermore, we highlight the advantages of computational reordering and emphasize the significance of memory/access order for modern GPUs.
Index Terms:
Compiler, Code Generation, GPUs, Program Optimization, Directive-Based ProgrammingI Introduction
@@
Over the last decade, the increasing popularity of computational accelerators in high-performance computing (HPC) has significantly impacted the programming landscape. The emergence of complex, unique, yet highly efficient architectures has posed a challenge for scientists, who must exert additional effort to take advantage of these architectures. Among the fastest computers, Graphics Processing Units (GPUs) are the most pervasive accelerators [1] and are inherently parallel, requiring software to expose concurrent calculations. Traditional optimization techniques may not provide performance benefits in such scenarios. Furthermore, with the approach of the end of Moore’s Law [2], accelerators have become increasingly domain-specific, and compilers have become essential for the automatic utilization of supercomputing resources.
@@
Although coding with low-level languages is typically the most profitable approach, utilizing them often involves complex data dependencies. As a result, adapting vendor/specific languages like CUDA [3] and OpenCL [4] requires extensive knowledge of the application and creates separate code entities just for computation offloading, which increases programming costs. To address this challenge, current GPU accelerators support programming with abstract models. OpenACC [5] and OpenMP [6] have emerged as popular models that alleviate the difficulty of accelerator use by providing code directives to existing languages. These directives allow users to specify compute/intensive parts of the original code, which compilers may automatically translate into accelerator code. At runtime, offloading is realized without any user intervention.
@@
Directive-based code typically maintains a general programming style and remains unspecialized until compilation. It is the responsibility of compilers to handle the computation, the redundancy of memory accesses, and the order of those while considering the nature of accelerators. However, code rewriting while preserving semantics may be challenging. Current compilers apply optimization for each metric sequentially, which limits opportunities for improved performance. In particular, user-specified directives fix loop structures, so little work has been done on optimising code under directives other than relying on compiler techniques used for sequential programs [7, 8]. As a result, more complex approaches are needed to improve the performance of directive-based code.
@@
This paper proposes the use of a modern optimization technique, equality saturation [9], to fine-tune directive/based GPU code. Our tool, ACC Saturator444The artifact is available at https://github.com/khaki3/acc-saturator., achieves less computation, fewer memory accesses, and high throughput at the same time, while easily integrating into the compilation of both OpenACC and OpenMP compilers. ACC Saturator performs kernel optimization by passing programs through an e-graph, a graph structure for equality saturation. Unlike other optimization methods, our approach neither transforms the abstract syntax trees nor it changes directives. Despite this, we attain significant performance improvements of up to 2.23x with the NVHPC compiler and 5.08x with GCC. We present a detailed performance analysis that highlights the benefits of each optimization on the state-of-the-art GPU architecture, NVIDIA Tesla A100. Our contributions are four-fold:
1.
@@
The rest of the paper is organized as follows. Section II provides background knowledge on GPUs, directive-based programming, code optimization, and equality saturation. Section III presents a high-level overview of our work. Sections IV and V explain our approach to represent general sequential code for equality saturation and optimizations. Section VI describes our novel code generation technique from e-graphs towards high/throughput GPU execution. In Section VII, we describe our experimental methodology. Section VIII presents the results of our evaluation along with in-depth analysis on each technique. Section IX discusses related work. Section X provides concluding remarks.
II Background
@@ II-A GPUs
Parallel computing is an assemblage of sequential execution, with means of intercommunication and mutual effects on efficiency. Such architectures accommodate simultaneous operations on multiple different data in uniform processors that are adjacent to each other. As an instance, a Graphics Processing Unit (GPU) runs hundreds of thousands of threads at the same time on around a hundred of streaming multiprocessors (SM). One of the most recent GPUs, NVIDIA Tesla A100, features 108 SMs and each SM contains 32 FP64 cores along with multiple LD/ST units [10].
@@
The primary benefit of GPU execution stems from overlapping utilization of both computational and memory units. Although GPUs provide high memory bandwidth, the global memory access latency is also higher than that of CPUs [11, 12, 13]. Therefore, optimal throughput may be attained by covering memory requests with computational execution and hiding the latency of data movement. Additionally, the order of memory accesses is essential due to the memory hierarchy, because the distance among accesses by neighboring threads often results in limited bandwidth [14].
II-B Directive-Based Programming
@@
Decomposing sequential execution into parallel entities is challenging. Routine calls often create data dependencies across multiple files, hindering the extraction of compute-intensive code. Even with extracted code, loops often possess intricate access patterns on different structures that challenge automatic parallelization for efficiency. Directive-based programming models, such as OpenACC [5] and OpenMP [6], alleviate the burden of introducing additional code for hardware acceleration. Both OpenACC and OpenMP provide code directives on the de facto standard languages for scientific applications (C/C++/ Fortran) to specify offloading parts as compute kernels with the declaration of explicit parallelism, enabling seamless automatic accelerator use.
@@
Listing 1 provides an example of OpenACC code that utilizes code directives to specify the parallelism of loops. The kernels directive guards a sequence of kernels, while the loop directive explicitly sets the parallelism of a loop. Users may also specify a single kernel region by using the parallel directive (not shown in the listing). OpenACC’s parallelism consists of three levels of abstraction: gang, worker and vector (in coarse to fine order). For instance, on the NVHPC compiler, both the top i and the middle j loops feature gang parallelism with the degree of the trip count cy and 16, respectively, and these are distributed over thread-blocks on GPUs. The middle j loop exposes additional vector parallelism and launches 256 threads to execute the inner statements. The blue comments show an equivalent form in OpenMP, which does not allow the reuse of parallelism across nested loops. Loop parallelism and data transfers may be implicit, and the compiler automatically solves data dependencies and sets optimal parameters. The NVHPC compiler generates embarrassingly parallel code for GPUs from directive code. On the other hand, GCC utilizes a principal-agent model for both OpenACC and OpenMP, and Clang follows the same approach for OpenMP [15].
II-C Limits of Annotation Optimization
GPU architectures require substantial parallelism to achieve high efficiency. Therefore, configuring parallelism appropriately based on the execution model is crucial. In directive-based programming, parallelism can be controlled through clauses like ’gang’, ’worker’, and ’vector’. These annotations significantly influence the mapping of code to hardware. For instance, the work by Lambert et al. [16]. illustrates that vector parallelism can lead to significant performance degradation by underutilizing thread-blocks. Conversely, Gong et al. [17]. demonstrate that specifying gang and vector parallelisms tailored to each compiler can lead to performance improvements over the compiler’s default settings. This suggests that application developers should fine-tune these directives based on actual hardware and software environments.
Moreover, once the parallelism annotations are set, any structural changes to the application may introduce performance bottlenecks. Techniques such as loop unrolling increase register usage, while loop folding may constrain parallelism, both of which reduce SM occupancy. Consequently, compilers must develop optimization strategies that enhance application performance not by altering the structural integrity but by refining the computational sequence. Our proposed ACC Saturator is designed to effectively accomplish this.
II-D Equality Saturation
@@
Compiler techniques have been critical in utilizing the best resources available for program execution on the target architecture for the last half-century [18, 19]. Various features of hardware, including registers, cache systems, pipelines, and parallelization, rely on the compiler’s efforts for performance, enabling target-specific optimization under resource constraints. However, compilers tend to miss a holistic view of performance opportunities and require additional discrete efforts for each discrete optimization mechanism [20]. Equality saturation, a state/of/the/art technique for compilers [9], defines a set of rewriting rules over the accumulation of equal expressions of target code. The rewriting continues until the expression gets saturated, finding no other forms, or hits a time or size limitation, attaining optimal codes based on a cost model that considers the entire computation.
@@
A dedicated graph structure called e-graph accepts the accumulation of equal expressions, while sharing redundant code blocks over expressions. The graph structure consists of e-classes, groups of equal e-nodes. One e-node may point to e-classes as its subexpressions, and the accumulation works by extending or merging e-classes in accordance with rewriting rules, while preserving the relationship between parental e-nodes and child e-classes [21, 22]. To extract optimal solutions, each e-class selects one contained e-node with a minimum cost, while optionally counting common expressions only once and completing common subexpression elimination (CSE).
III Methodology Overview
@@
The e-graph operation by ACC Saturator consists of three phases. First, we build a static single-assignment form (SSA) from the input code to create initial e-graphs (Section IV). E-graphs accept our SSA representation holding dependencies, conditions, and iterations over C-style operations. Second, we define rewriting rules and a cost model for performance improvement (Section V). Our tool runs equality saturation and extracts optimal solutions under time and size limitations. Last, we generate output code (Section VI). As a result, user kernels are reduced to minimum computation with a new order.
IV Program Representation
During compiler optimizations, the program semantics and behavior are preserved while improving its performance. Maintaining the order of data accesses is especially important to ensure reproducibility of results. To express clear dependencies among statements, program analysis often relies on static single-assignment form (SSA), which allows only one definition per variable [23]. Early work on equality saturation [9] used an SSA-based graph structure to represent and rewrite an entire program. For directive/based code, however, user-specified parallelism significantly affects performance [24], so compilers are limited to respect users’ decisions.
@@ IV Program Representation
ACC Saturator is a more practical compiler optimization framework than previous work [9] because it maintains code structures, is capable of restoring the original state, and is applicable to standard languages such as C and Fortran. ACC Saturator allows equality saturation over directive-based code, upholding the importance of structure information which previous work disregarded. Reusing the code style that the input provides, ACC Saturator can support the user’s intended target architecture by rewriting rules that lead to a new order of computation. Each rule is simple, but by combining many rules, we discover better solutions and attain speedups over the state-of-the-art industry compilers.
IV-A E-Graph Creation
@@
Both OpenACC and OpenMP contain sequential parts within the innermost parallel loops that may be executed independently across multiple threads and thus optimized for more efficient execution or reduced computation. To attain this, we create an e-graph for each innermost parallel loop as follows: First, we introduce conditional nodes [25] to represent control structures such as if and for, while merging data flows. Second, we assign an ID to each variable/array assignment or . Third, we update each variable/array load to refer to the latest ID along its data flow. Last, for each assignment or , we assign both the ID and the expression to the same e-class.
Examples of e-graphs for store/load, if, and for operations are illustrated in at Figure 1, with some IDs omitted for simplicity. The e-nodes are depicted as white blocks, and the e-classes as dotted boxes, while the black arrows represent parent-child relationships among them. As shown in each example, multiple references to the same variables are now directed toward common e-classes, thereby reducing redundancy. For each updated variable, is created within if and for structures with conditions that may be concrete or abstract. Our tool supports C-style operations, such as function calls, pointers, and member references, within the same SSA framework. To optimize the code, ACC Saturator leverages our e-graph representation with rewriting rules and cost models, as detailed in Section V.
IV-B Code Selection
@@
To select optimal codes in the e-graph, we aim to find an equal expression that corresponds to a set of all assignments. We extract the lowest-cost expression that contains all the e-classes of assignments based on a sequence of IDs. The total cost is calculated as the sum of the cost of each e-class, with common e-classes being counted only once. To attain this, we use linear programming techniques [26].
@@
Each assignment is updated based on its ID and the extracted expression. To enable reuse across assignments, the values of common expressions are stored in temporary variables. ACC Saturator allows for customizing the generation order of these temporary variables (Section VI).
V Optimization with Saturation
In this section, we outline our methodology for enhancing OpenACC/OpenMP code performance through equality saturation. Our approach is based on a meticulously developed set of rewriting rules paired with a cost model designed to optimize computation and memory access efficiencies. As highlighted in Section II-C, structural modifications in the code may cause significant performance variations. Additionally, the architectural specifics in directive-based programming often remain undisclosed. To address these challenges, our primary strategy involves optimizing the sequence and volume of computations, providing a generalized solution to accelerate application performance.
V-A Rewriting Rules
Reordering computations allows programs to explore different optimization possibilities. Rather than selecting a single alternative, compilers aim to choose the most profitable candidate according to their objective. Equality saturation defers this decision and instead accumulates equal expressions using multiple rewriting rules. With this approach, ACC Saturator can derive efficient operations and facilitate the reuse of expressions simultaneously, as shown in at Figure 1.
@@ V-A Rewriting Rules
We apply two sets of rewriting rules to our equality saturation process. The first set introduces fused multiply/adds (FMA) operations, which can improve code generation and resource utilization on GPUs [20]. Table I lists the minimum set of rules we use. When we encounter expressions that match the FMA pattern, we add the corresponding operations to the e-classes of matched expressions. The second set of rules benefits from the commutative and associative properties of the plus and multiply operators to reorder computation. This can enable common subexpression elimination and produce new FMA operations. We also incorporate constant folding of arithmetic operations with integer and floating-point numbers.
@@
V-B Cost Model
GPUs are complex systems that require careful analysis to accurately predict their efficiency [27, 28]. While applications running on GPUs often face memory-bound limitations [29], it is not just the number of memory accesses that affects overall bandwidth. Factors such as on-chip resource utilization, processor occupancy, thread/grid-level parallelism across multiple memory layers, and instruction-level parallelism (ILP) all play a role, and improving one metric often comes at the expense of another. In ACC Saturator, our focus is on reducing memory access and computation by utilizing registers for common expressions, while maintaining ILP as described in Section VI.
@@ VI Code Generation
VI-A Temporary-Variable Insertion
at Figure 1 depicts the extracted expression in a directed graph. For each selected e-node, ACC Saturator generates a temporary variable to store the computational result, placing it immediately before the corresponding use. In cases where multiple statements reference an e-node, we select the innermost scope to declare a variable for those statements. Every assignment modifies the right-hand expression to include the variable of the corresponding e-node.